# Active learning emulators for nuclear two-body scattering in momentum space

### v1

@@ IV Summary and Outlook
In this work, we have extended the active learning emulators G-ROM and LSPG-ROM, developed by Maldonado et al. [Maldonado:2025ftg] for two-body scattering with error estimation in coordinate space, to coupled-channel scattering in momentum space. Our full-order model (FOM) solver is based on the Lippmann-Schwinger integral equation for the scattering $t$-matrix as opposed to the radial Schrödinger equation considered in Ref. [Maldonado:2025ftg]. Hence, our approach gives access to a much wider range of modern chiral NN interactions, including the high-order SMS potentials developed by LENPIC [Reinert:2017usi, Epelbaum:2022cyo].
Specifically, our emulators are based on Galerkin (G-ROM) and Petrov-Galerkin (LSPG-ROM) projections, as well as on high-fidelity calculations at a few snapshots across the parameter space of the chiral NN interactions [Melendez:2022kid, Drischler:2022ipa]. Chiral NN interactions exhibit an affine dependence on the short-range contact interactions, which we considered here as parameters, enabling efficient offline-online decompositions. A greedy algorithm estimates the emulator error and iteratively enriches the emulator basis by adding snapshots to regions where the estimated error is largest, until a specified error tolerance is met [hesthaven2015certified, Quarteroni:218966, Benner20201]. Similarly to Ref. [Maldonado:2025ftg], we found that this active learning approach to emulator training has two significant advantages over the naïve POD approach: It typically requires much fewer (computationally demanding) high-fidelity calculations to achieve similar accuracy, and includes emulator error estimation. Both features are critical for conducting principled uncertainty quantification of scattering.
Overall, we have found that the greedy algorithm performs similarly well in momentum space for constructing ROMs as in coordinate space [Maldonado:2025ftg]. For example, it converges exponentially when applied to the GT+ potential in the uncoupled $^{1}\text{S}_{0}$ and the coupled $^{3}\text{S}_{1}$$^{3}\text{D}_{1}$ channels, with accuracies comparable to those of the POD approach. After this convergence analysis, we presented results for emulating phase shifts in the deuteron channel (as a representative channel) and total cross sections. The latter calculations involve emulating phase shifts across various (coupled and uncoupled) partial-wave channels. In all cases, the emulators reproduced the high-fidelity calculations with high accuracy.
@@
Emulator speed-up factors depend inherently on the software and hardware implementations. To assess the computational speed-up factors more rigorously, without being overly exhaustive, we implemented highly efficient FOMs and ROMs in Python using JAX. We have found that our emulators can be $100$ times faster (or even more) than the already efficient FOM solver, depending on the emulator’s basis size, enabling efficient Bayesian parameter estimation, even at the two-body level.
@@
Our emulators, together with our recent advances in developing active learning emulators for 3N scattering [Gnech:2025gsy, Gnech:2025lbg], constitute a significant step toward comprehensive Bayesian calibrations of chiral NN and 3N forces using scattering data. These Bayesian analyses will be critical for elucidating issues in chiral EFT [Furnstahl:2021rfk], including the efficacy of modified power countings relative to the popular Weinberg power counting and the differing predictions of chiral EFT for ground-state energies and charge radii of atomic nuclei in the medium- to heavy-mass regime.
@@
Future work will focus on combining the developed NN and 3N emulators into a unified framework for systematic Bayesian analysis of scattering observables. To achieve this goal, several developments are necessary. Our emulators with error estimation need to be extended to other scattering observables, such as differential cross sections and spin observables. Furthermore, to emulate non-affine model parameters, such as the momentum cutoff and pion mass, versatile hyperreduction methods, including rational function approximations, need to be developed to enable efficient emulator offline-online decompositions. At the NN level, the empirical interpolation method (EIM) has already been tested and demonstrated to be efficient at rendering Woods-Saxon-like potentials approximately affine [Odell:2023cun, Catacora-Rios:2025pau].
@@
In addition, accurate and efficient methods for estimating the minimum singular values of the (asymmetric) FOM matrix $\mathbf{A}({\bf\it\theta})$ in the parameter space need to be developed to derive more rigorous upper bounds on the emulator errors (see also Ref. [Maldonado:2025ftg]). These would enable conservative error estimation. Furthermore, the 3N emulators developed in Refs. [Gnech:2025gsy, Gnech:2025lbg] need to be extended to energies above the deuteron breakup threshold, a project that is currently under development by the STREAMLINE2 collaboration.
@@
The ultimate goal is to apply this unified framework for emulating scattering observables to calibrate next-generation chiral NN and 3N forces, with emulator errors and EFT truncation errors fully quantified using the BUQEYE machinery [Melendez:2017phj, Melendez:2019izc, Wesolowski:2021cni, Svensson:2022kkj, Svensson:2023twt], and study their predictions for atomic nuclei and infinite matter. These developments will benefit from the software framework we will make publicly available, thereby enabling broader adoption of active learning emulators with error estimation within the nuclear physics community.
@@ Acknowledgements.
We thank Petar Mlinarić, Stephan Rave, and our STREAMLINE2 collaborators for fruitful discussions, and Ingo Tews for sharing the source code that evaluates the chiral potentials developed in Ref. [Gezerlis:2014zia]. We also thank ECT* for support at the 2025 workshop “Next generation ab initio nuclear theory” during which this work was presented and further developed. This material is based upon work supported by the U.S. Department of Energy, Office of Science, Office of Nuclear Physics under contract No. DE-FG02-93ER40756, under the FRIB Theory Alliance award DE-SC0013617, under the STREAMLINE collaboration awards DE-SC0024233 (Ohio University) and DE-SC0024509 (Ohio State University), the STREAMLINE2 collaboration award DE-SC0026198, and by the National Science Foundation under awards PHY-2209442/PHY-2514765. Computational resources for this work were provided by the NSF MRI grant 2320493 and the National Energy Research Scientific Computer Center (NERSC), a U.S. Department of Energy Office of Science User Facility supported by the Office of Science of the U.S. Department of Energy under Contract No. DE-AC02-05CH11231 using NERSC award NP-ERCAP0033451. The following open-source libraries were used to generate the results in this work: corner [corner], cython [behnel2011cython], emcee [Foreman_Mackey:2013aa], JAX [jax2018github], GSL [gsl], matplotlib [Hunter:2007], mpmath [mpmath], numpy [harris2020array], and scipy [2020SciPy-NMeth].
@@ A.1 Single channel equation
which can be solved by standard linear algebra routines. For the high-fidelity solutions, we use the SciPy’s wrapper around LAPACK’s efficient solver ZGESV for general matrices, linalg.solve() [2020SciPy-NMeth].
@@ Appendix B Calculation of phase shifts, inelasticity parameters, and total cross sections
In this appendix, we present the calculation of phase-shifts, inelasticity parameters, and total cross sections from the on-shell $t$-matrix elements. For completeness, we also present the explicit diagonalization of the coupled-channels according to the Stapp [Stapp:1956mz] parameterization. We conclude with the expression for the total cross section.
@@
For the coupled channel states, the partial wave $S$-matrix elements can be parameterized in the Stapp et al. [Stapp:1956mz] form, using the “bar” phase shifts $\bar{\delta}_{-}$, $\bar{\delta}_{+}$ and the mixing parameter $\bar{\varepsilon}$ by means of the transformation
@@
The calculations of the high-fidelity solutions in the present work have been validated by comparing them to the codes of Refs. [Elster:1988pp, Elster:1988zu].
Using the bar phase shifts and inelasticities, we can define amplitudes $\tau^{j}_{\pm}(E_{k_{0}})$ similar to Eq. (69). We calculate the total cross sections using the optical theorem. In terms of scattering amplitudes $\tau^{j}$ it can be written as [Silbar:1979wp]

### v2

@@ IV Summary and Outlook
In this work, we have extended the active learning emulators G-ROM and LSPG-ROM, developed by Maldonado et al. [28] for two-body scattering with error estimation in coordinate space, to coupled-channel scattering in momentum space. Our FOM solver is based on the Lippmann-Schwinger integral equation for the scattering $t$-matrix as opposed to the radial Schrödinger equation. Hence, our approach gives access to a much wider range of modern chiral NN interactions, including the high-order SMS potentials developed by LENPIC [40, 13].
Specifically, our emulators are based on Galerkin (G-ROM) and Petrov-Galerkin (LSPG-ROM) projections, as well as on high-fidelity calculations at a few snapshots across the parameter space of the chiral NN interactions [30, 7]. Chiral NN interactions exhibit an affine dependence on the short-range contact interactions, which we considered here as parameters, enabling efficient offline-online decompositions. A greedy algorithm estimates the emulator error and iteratively enriches the emulator basis by adding snapshots to regions where the estimated error is largest, until a specified error tolerance is met [22, 39, 1]. Similarly to Ref. [28], we found that this active learning approach to emulator training has two significant advantages over the naïve POD approach: It typically requires much fewer (computationally demanding) high-fidelity calculations to achieve similar accuracy, and includes emulator error estimation. Both features are critical for conducting principled uncertainty quantification of scattering.
Overall, we have found that the greedy algorithm performs similarly well in momentum space for constructing ROMs as in coordinate space [28]. For example, it converges exponentially when applied to the GT+ potential in the uncoupled $^{1}\text{S}_{0}$ and the coupled $^{3}\text{S}_{1}$$^{3}\text{D}_{1}$ channels, with accuracies comparable to those of the POD approach. After this convergence analysis, we presented results for emulating phase shifts in the deuteron channel (as a representative channel) and total cross sections. The latter calculations involve emulating phase shifts across various (coupled and uncoupled) partial-wave channels. In all cases, the emulators reproduced the high-fidelity calculations with high accuracy.
@@
Emulator speedup factors depend inherently on the software and hardware implementations. To assess the computational speedup factors more rigorously, without being overly exhaustive, we implemented highly efficient FOMs and ROMs in Python using JAX. We have found that our emulators can be $100$ times faster (or even more) than the already efficient FOM solver, depending on the emulator’s basis size, enabling efficient Bayesian parameter estimation, even at the two-body level.
@@
Our emulators, together with our recent advances in developing active learning emulators for 3N scattering [20, 21], constitute a significant step toward comprehensive Bayesian calibrations of chiral NN and 3N forces using scattering data. These Bayesian analyses will be critical for elucidating issues in chiral EFT [16], including the efficacy of modified power countings relative to the popular Weinberg power counting and the differing predictions of chiral EFT for ground-state energies and charge radii of atomic nuclei in the medium- to heavy-mass regime.
@@
Future work will focus on combining the developed NN and 3N emulators into a unified framework for systematic Bayesian analysis of scattering observables. To achieve this goal, several developments are necessary. Our emulators with error estimation need to be extended to other scattering observables, such as differential cross sections and spin observables. Furthermore, to emulate non-affine model parameters, such as the momentum cutoff and pion mass, versatile hyperreduction methods, including rational function approximations, need to be developed to enable efficient emulator offline-online decompositions. At the NN level, the empirical interpolation method (EIM) has already been tested and demonstrated to be efficient at rendering Woods-Saxon-like potentials approximately affine [37, 5].
@@
In addition, accurate and efficient methods for estimating the minimum singular values of the (asymmetric) FOM matrix $\mathbf{A}({\bf\it\theta})$ in the parameter space need to be developed to derive more rigorous upper bounds on the emulator errors (see also Ref. [28]). These would enable conservative error estimation. Furthermore, the 3N emulators developed in Refs. [20, 21] need to be extended to energies above the deuteron breakup threshold, a project that is currently under development by the STREAMLINE2 collaboration.
@@
The ultimate goal is to apply this unified framework for emulating scattering observables to calibrate next-generation chiral NN and 3N forces, with emulator errors and EFT truncation errors fully quantified using the BUQEYE machinery [33, 32, 49, 45, 46], and study their predictions for atomic nuclei and infinite matter. These developments will benefit from the software framework we made publicly available [4], thereby enabling broader adoption of active learning emulators with error estimation within the nuclear physics community.
@@ Acknowledgements.
We thank Petar Mlinarić, Stephan Rave, and our STREAMLINE2 collaborators for fruitful discussions, and Ingo Tews for sharing the source code that evaluates the chiral potentials developed in Ref. [18]. We also thank ECT* for support at the 2025 workshop “Next generation ab initio nuclear theory” during which this work was presented and further developed. This material is based upon work supported by the U.S. Department of Energy, Office of Science, Office of Nuclear Physics under contract No. DE-FG02-93ER40756, under the FRIB Theory Alliance award DE-SC0013617, under the STREAMLINE collaboration awards DE-SC0024233 (Ohio University) and DE-SC0024509 (Ohio State University), the STREAMLINE2 collaboration award DE-SC0026198, and by the National Science Foundation under awards PHY-2209442/PHY-2514765. Computational resources for this work were provided by the NSF MRI grant 2320493 and the National Energy Research Scientific Computer Center (NERSC), a U.S. Department of Energy Office of Science User Facility supported by the Office of Science of the U.S. Department of Energy under Contract No. DE-AC02-05CH11231 using NERSC award NP-ERCAP0033451. The following open-source libraries were used to generate the results in this work: corner [corner], cython [behnel2011cython], emcee [Foreman_Mackey:2013aa], JAX [3], GSL [gsl], matplotlib [Hunter:2007], mpmath [mpmath], numpy [harris2020array], and scipy [48].
@@ A.1 Single channel equation
which can be solved by standard linear algebra routines. For the high-fidelity solutions, we use the SciPy’s wrapper around LAPACK’s efficient solver ZGESV for general matrices, linalg.solve() [48].
@@ Appendix B Calculation of phase shifts, inelasticity parameters, and total cross sections
In this appendix, we present the calculation of phase-shifts, inelasticity parameters, and total cross sections from the on-shell $t$-matrix elements. For completeness, we also present the explicit diagonalization of the coupled-channels according to the Stapp [43] parameterization. We conclude with the expression for the total cross section.
@@
For the coupled channel states, the partial wave $S$-matrix elements can be parameterized in the Stapp et al. [43] form, using the “bar” phase shifts $\bar{\delta}_{-}$, $\bar{\delta}_{+}$ and the mixing parameter $\bar{\varepsilon}$ by means of the transformation
@@
The calculations of the high-fidelity solutions in the present work have been validated by comparing them to the codes of Refs. [11, 12].
Using the bar phase shifts and inelasticities, we can define amplitudes $\tau^{j}_{\pm}(E_{k_{0}})$ similar to Eq. (69). We calculate the total cross sections using the optical theorem. In terms of scattering amplitudes $\tau^{j}$ it can be written as [42]